# Non-monotonic emergence of order from chaos in turbulent thermo-acoustic fluid systems

### v3

@@ I Introduction
We perform experiments in a turbulent thermo-acoustic fluid system (Fig. 1(a)), which is known to exhibit dynamical transition from chaos to order. Thermo-acoustic systems find applications in gas turbine combustors and rocket engines. The turbulence intensity in a closed duct increases with increasing Reynolds number $Re$ [19]. Turbulence introduces a wide range of time and length scales [20], thus promoting chaotic dynamics. Also, flow in a confinement creates an acoustic field. As the flow rate increases, the system exhibits a transition from low-amplitude aperiodic to high-amplitude periodic fluctuations in the acoustic pressure signal. We note that even though the acoustic pressure exhibits self-sustained limit cycle oscillations, the underlying flow and flame dynamics are still turbulent. Such self-sustained oscillations emerge due to to interactions between the flow, flame, and the acoustic subsystem [21, 22, 23].
Fluctuations in one subsystem depend on the function of other subsystems and the spatio-temporal structure of the inter-subsystem interactions [23]. While turbulence promotes disorder and chaos, nonlinear feedback between subsystems promotes self-organization and order. Although the range of spatial and temporal scales increases with $Re$, we observe the emergence of a single dominant frequency in acoustic pressure, periodic shedding of large coherent structures, and sudden and intense heat release patterns occurring periodically [24]. Since the subsystems are co-evolving, the fluctuations in each subsystem determine as well as necessarily reflect the evolution of interactions between all subsystems. Thus, it is essential to study the role and characteristics of fluctuations during dynamical transitions.
Turbulent thermo-acoustic fluid systems are known to exhibit a transition in the acoustic field from low-amplitude aperiodic fluctuations (which is deterministic chaos [25]) as shown in Fig. 1(d), to high-amplitude periodic oscillations (order) as shown in Fig. 1(f) with an increase in the Reynolds number ($Re$) of the inlet flow [22, 13]. Such a transition occurs via the route of intermittency [26], which is characterized by bursts of high-amplitude periodic oscillations amidst epochs of low-amplitude chaotic oscillations in the acoustic field (Fig. 1(e)). The ordered dynamics obtained at high $Re$ is a high-amplitude limit cycle oscillation, termed thermo-acoustic instability, which can cause catastrophic destruction to engines and restrict their operational envelope [22, 13].
@@
Here, we use complex networks to encode the fluctuations in the temporal dynamics of acoustic pressure fluctuations ($p^{\prime}$). A network is derived from time series of $p^{\prime}$ obtained from experiments in a turbulent combustor, where amplitude bins are nodes, and the weight of a link connecting two nodes defines the probability of transition between the respective amplitude bins [37, 38, 39]. We refer to the resulting network as amplitude transition network. We construct these networks at distinct timescales and examine the network properties. Using network entropy, we discover that the uncertainty and number of amplitude transitions first increase and then decrease during the transition from chaos to order in the system, delineating a non-monotonic emergence of ordered dynamics from disorder. We also characterize the amplitude scales in the fluctuations during the order emergence using distance-based centrality measures. The structure of the article is as follows: The details of the experimental setup and data acquisition are described in Sec. II.1. The network construction and the choice of parameters involved are explained in Sec. II.2. The measures derived from the network and their interpretations are described in Sec. III. Our conclusions are given in Sec. IV.
@@ II.1 Experimental setup
We conducted experiments in a turbulent combustor with two different types of flame-holding devices. A schematic of the turbulent combustor used in this study is shown in Fig. 1(a). The turbulent combustor comprises a plenum chamber, a burner, and a combustion chamber with extension duct. Another chamber with a relatively larger cross-sectional area, known as the decoupler, is connected to one end of the combustion duct. The decoupler ensures that the pressure fluctuations at the end of the duct is the same as the ambient air. A central shaft of diameter 16 mm through the burner supports the flame-stabilizing circular disc object, i.e., a bluff-body (Fig. 1(b)) or a vane swirler (Fig. 1(c)).
@@
Compressed air enters the plenum chamber of the combustor before passing through the burner, where the mixing of the air and fuel (liquefied petroleum gas) takes place. The central shaft delivers fuel into the combustion chamber via four radial injection holes with diameters of 1.7 mm each. The dynamics of the unsteady acoustic pressure oscillations in the combustor is investigated as the system parameter, i.e., the Reynolds number of the incoming flow is varied. The acoustic pressure fluctuations in the combustion chamber are acquired using a piezoelectric pressure transducer (PCB103B02, with an uncertainty of $\pm$ 0.15 Pa). The transducer is mounted at a distance of 25 mm downstream from the dump plane and records the pressure signal for 3 s at a sampling rate of 10 kHz.
@@
The $Re$ near the position of the burner is calculated as $Re=4\dot{m}D_{1}/\pi\mu D_{0}^{2}$, where $\dot{m}$ is the sum of air-flow rate ($\dot{m_{a}}$) and fuel-flow rate ($\dot{m_{f}}$), $D_{0}$ is the hydraulic diameter of the burner, $D_{1}$ is the diameter of the circular bluff-body ($D_{1}=D_{0}$ for experiments with swirler) and $\mu$ is the kinematic viscosity of the fuel-air mixture at the experimental condition. We use the expression for dynamic viscosity as stated by Wilke [40] to compute the $Re$ of a binary gas mixture.
@@
The fuel-air mixture is injected at turbulent conditions and undergoes mixing in recirculation zones created by a flame stabilizing mechanism. Here, we perform two sets of experiments using two types of flame stabilizers, namely, a bluff body and a vane swirler. The bluff body located at the end of the shaft forces the flow to turn near its shoulder, creating a large-scale recirculation zone downstream of the bluff body. On the other hand, a swirler is a non-rotating device with curved vanes (blade-like surfaces) placed radially around a central shaft, forcing the flow to turn along its vanes, creating a large-scale recirculation zone downstream of the swirler. This zone comprises hot radicals, and the large-scale recirculation vortices enhance the fuel-air mixing, which subsequently undergoes combustion. The heat released due to the combustion of the fuel-air mixture adds energy to the fluctuations in the acoustic field, leading to growth in its amplitude. Perturbations in the acoustic field influence the flame surface oscillations, leading to fluctuations in the heat release rate. Further, the acoustic and hydrodynamic fluctuations together influence the mixing and convection of the fuel-air mixture via coherent structures in the flow, which subsequently determine the dynamics of the turbulent thermo-fluid system [21, 22].
@@
For the experiments on the turbulent combustor with a bluff-body flame stabilizer, the fuel flow is set constant at 48 SLPM (standard liter per minute), and the airflow is varied from 752 – 1446 SLPM quasi-statically and hence the Reynolds number is varied from $(3.865\pm 0.051)\times 10^{4}$ to $(6.759\pm 0.072)\times 10^{4}$. Similarly, for the experiments on the turbulent combustor with vane swirl flame stabilizer, we set a constant fuel flow rate of 21 SLPM, and the airflow rate is varied from 330 – 550 SLPM quasi-statically (i.e., the Reynolds number is varied from $(1.441\pm 0.020)\times 10^{4}$ to $(2.752\pm 0.032)\times 10^{4}$.
@@
II.2 Construction of amplitude transition networks
We use the stochastic mapping network construction technique proposed by Shirazi et al. [37], which explicitly encodes the temporal sequence of the considered variable. For a discrete time series $x(t)$ = $\{x_{1},x_{2},\ldots,x_{n}\}$, we begin the network construction by grouping these data points into horizontal amplitude-bins, i.e., discrete ‘states.’ The number of bins required and the respective width of each bin are chosen such that the number of data points in each bin is approximately equal. The number of bins ($N$) needed for this formulation is given by $N\approx 2n^{2/5}$ [43].
@@ II.3 Measures from amplitude transition network
We construct amplitude-transition networks [37, 38] using the normalized time series of acoustic pressure fluctuations ($p^{\prime}$). The normalization is done with the maximum value of the acoustic pressure during the state of thermo-acoustic instability. As described in Sec. II.2, a weighted and directed network is constructed for each dynamical state at different time scales $\tau$. The topology of the constructed network encodes the temporal order of fluctuations in the dynamics. Periodic oscillations would result in a regular network topology, where connections exist between nodes that correspond to adjacent amplitude bins. On the other hand, chaotic fluctuations would be encoded as connections between nodes corresponding to non-adjacent amplitude bins and corresponding to different amplitude scales, and the network structure becomes more disorderly.
@@
In amplitude transition network, at a specific timescale, the number of possible transitions from one amplitude scale to another is infinitely many. Yet, only a selected multitude of these transitions are realized during a certain dynamical state, reflecting the ‘complexity’ of the fluctuation dynamics at the selected timescale. Complexity here simply refers to the number of possible realizations or transitions a variable can exhibit at a fixed timescale [49, 50]. For example, a periodic signal has least complexity as only certain amplitude transitions are realized in an orderly manner in the system. Whereas, during chaos, many amplitude transitions are realized, and the occurrence of any one of these transitions is highly uncertain. Hence, we emphasize that entropy is a measure of complexity only at the timescale at which the amplitude transitions are encoded in the network [49]. Also, Shannon entropy is equivalent to the Clausius entropy, which characterizes the macroscopic thermodynamic property of a system [51]. Thus, $\mu_{S}$ of an amplitude-transition network encodes the uncertainty in temporal fluctuations of a dynamical system and also quantifies the thermodynamic disorder of the system.
Figure 3: The average Shannon entropy of transition networks, which quantifies the amount of disorder, first increases then decreases during the transition from chaos to order for different transition time-scale $\tau$ (in $10^{-4}$ s). (a) Variation of average network entropy ($\mu_{S}$) and root-mean-square of the non-dimensional pressure fluctuations $p^{\prime}_{rms}$ with the flow control parameter $Re$ for a bluff-body stabilized turbulent combustor. The shaded region corresponds to the region where $Re<(Re_{c})_{\text{bluff}}$ where $(Re_{c})_{\text{bluff}}=4.966\times 10^{4}$. (b) Variation of average network entropy ($\mu_{S}$) and root-mean-square of the pressure fluctuations ($p^{\prime}_{rms}$) with the flow control parameter $Re$ for a swirl stabilized turbulent combustor. The shaded region corresponds to the region where $Re<(Re_{c})_{\text{swirl}}$ where $(Re_{c})_{\text{swirl}}=1.724\times 10^{4}$.
Figure 4: (a) Variation of average information entropy ($\mu_{S}$) of the network constructed from the time series corresponding to the dominant acoustic mode at different transition time scale $\tau$ (in $10^{-4}$ s). We chose the optimal time-scale to be one that maximizes the information obtained from the network, i.e., $\tau_{opt}=14\times 10^{-4}$ s for both bluff-body and swirl stabilized turbulent combustor. (b) $\mu_{S}$ from both bluff-body and swirl combustor corresponding to $\tau_{opt}$ exhibits a power law scaling $(\mu_{S})_{\text{bluff}}\propto(p^{\prime}_{rms})^{\alpha_{1,2}}$, and $(\mu_{S})_{\text{swirl}}\propto(p^{\prime}_{rms})^{\beta_{1,2}}$, respectively with similar exponents. The power law exponents for a bluff-body combustor when $Re<(Re_{c})_{\text{bluff}}$ is $\alpha_{1}=0.09\pm 0.025$, and when $Re>(Re_{c})_{\text{bluff}}$ is $\alpha_{2}=-0.20\pm 0.009$. Similarly, the power law exponents for a swirl combustor when $Re<(Re_{c})_{\text{swirl}}$ is $\beta_{1}=0.15\pm 0.049$, and when $Re>(Re_{c})_{\text{swirl}}$ is $\beta_{2}=-0.21\pm 0.020$.
@@
Furthermore, we explore the diversity in the magnitude of jumps between different amplitude scales associated with fluctuations. To do so, we use path-based centrality measures derived from amplitude-transition networks. A path from one node to another in the network denotes the jump from one amplitude bin to another in the time series. Thus, we compute distance-based metrics, namely the characteristic path length (CPL) and mean betweenness centrality ($\overline{C}_{BC}$).
@@
The CPL of a network is the mean number of steps needed to reach one node from another in the network [52, 53]. For amplitude transition networks, CPL encodes the magnitude of jump in amplitude scale due to fluctuations between different amplitude bins (nodes). For example, in a periodic time series, any amplitude bin (node) is predominantly connected to its adjacent nodes. The differences in amplitude scales are small and of similar magnitude when the transition occurs from any one bin to its neighboring bin, and this is reflected as a high value of CPL. On the other hand, a small value of CPL indicates that fluctuations occur between nodes of very different amplitude scales, such as during chaotic dynamics.
@@
The betweenness centrality of a node quantifies the extent to which a node lies in the path between other nodes of the network [54, 53]. For an amplitude transition network, the mean betweenness centrality $\overline{C}_{BC}$ quantifies the order in which transition occurs between different amplitude scales. If there are direct transitions between bins of diverse amplitude scales, such as during chaotic fluctuations, then the value of $\overline{C}_{BC}$ will be low. However, for a periodic time series, the connections are mostly between adjacent nodes, and any node lies ‘in between’ the path of many other node pairs with different amplitude scales, resulting in a high value for $\overline{C}_{BC}$.
@@ III Results
Figure 5: Quantifying diversity in amplitude scale of fluctuations using distance-based network topological metrics. (a) Variation of characteristic path length (CPL) of network constructed from time series of non-dimensional $p^{\prime}$ of bluff-body and swirl stabilized turbulent combustor with $p^{\prime}_{rms}$ (i.e., for the entire range of control parameter $Re$) at $\tau_{opt}$. The measure from both combustors exhibits the same power law scaling $\text{CPL}\propto(p^{\prime}_{rms})^{\gamma}$, with exponent $\gamma=0.28\pm 0.016$. (b) Variation of average betweenness centrality ($\overline{C}_{BC}$) of network constructed from time series of non-dimensional $p^{\prime}$ of bluff-body and swirl stabilized turbulent combustor with $p^{\prime}_{rms}$ at $\tau_{opt}$. $\overline{C}_{BC}$ from the bluff-body and swirl combustor exhibits a similar power law scaling $\overline{C}_{BC}\propto(p^{\prime}_{rms})^{\kappa}$ with exponents $\kappa=0.48\pm 0.033$ and $\kappa=0.47\pm 0.036$, respectively.
@@
Figure 3 shows the variation in $\mu_{S}$ of networks derived at different transition time scales $\tau$ (in $10^{-4}$ s; as the sampling frequency of the time series is 10 kHz) from the acoustic pressure signal as the control parameter ($Re$) is varied quasi-statically in experiments. We find that the average network entropy $\mu_{S}$ varies non-monotonically with $Re$ during the transition from chaos to order in both bluff-body (Fig. 3(a)) and swirl stabilized combustors (Fig. 3(b)) in a similar manner for several transition time-scales $\tau$. $\mu_{S}$ increases up to a critical Reynolds number ($Re_{c}=4.966\times 10^{4}$ for bluff-body combustor and $Re_{c}=1.724\times 10^{4}$ for swirl combustor) and then decreases. In Appendix. A, we show that the trend in network entropy is largely unaffected with variation in the other network construction parameter, i.e., the number of amplitude bins ($N$).
@@
Next, we chose an optimal transition timescale ($\tau_{opt}$) as the value of $\tau$ for which $\mu_{S}$ of the amplitude transition network of $p^{\prime}$ during the state of limit cycle oscillations is maximum and/or saturates. We obtain $\tau_{opt}=14\times 10^{-4}$ s for both the bluff-body and the swirl combustor (Fig. 4(a)). The variation of $\mu_{S}$ with $p^{\prime}_{rms}$ at $\tau_{opt}$ is shown in Fig. 4(b). We find that $\mu_{S}$ exhibits a power law scaling with $p^{\prime}_{rms}$ and the power law exponents is similar for the results obtained from both set of experiments (bluff-body and swirler configuration). Such scaling relation is of particular interest as it helps relate the uncertainty in amplitude transitions to the amplitude itself. Evidently, this relation is similar for different combustors and indicates universality in the features of the dynamical transition rather than dependence on the specific configuration of the system.
@@
The non-monotonicity of entropy during the emergence of order reveals that during the transition from chaotic to periodic dynamics, there is an interplay between the mechanisms that promote order and disorder. A thermo-acoustic system is a complex system where the acoustic, combustion, and hydrodynamic subsystems are nonlinearly coupled, leading to synchronized co-evolution in these subsystems and promoting pattern formation and ordered dynamics. On the other hand, as we increase the Reynolds number of the inlet flow, the turbulent intensity in the confined duct increases, promoting disorder. Our findings imply that disorder created by increasing turbulence intensity is dominant until a critical control parameter $Re_{c}$. Beyond this $Re_{c}$, the nonlinear feedback dominates, causing disorder to decay monotonically. Moreover, the decrease in $\mu_{S}$ beyond $Re_{c}$ is a precursor for an impending thermo-acoustic instability.

### v4

@@ I Introduction
We perform experiments in a turbulent thermo-acoustic fluid system (Fig. 1(a)), which is known to exhibit dynamical transition from chaos to order. Thermo-acoustic systems find applications in gas turbine combustors and rocket engines. The turbulence intensity in a closed duct increases with increasing Reynolds number $Re$ [19]. Turbulence introduces a wide range of time and length scales [20, 21], thus promoting chaotic dynamics in the acoustic field. As the inlet air flow rate increases, the system exhibits a transition from low-amplitude aperiodic to high-amplitude periodic fluctuations in the acoustic pressure signal. We note that even though the acoustic pressure exhibits self-sustained limit cycle oscillations, the underlying flow and flame dynamics are still turbulent. Such self-sustained oscillations emerge due to interactions between the flow, flame, and the acoustic subsystem [22, 23, 24].
Fluctuations in one subsystem depend on the function of other subsystems and the spatio-temporal structure of the inter-subsystem interactions [24]. While turbulence promotes disorder and chaos, nonlinear feedback between subsystems promotes self-organization and order. Although the range of spatial and temporal scales increases with $Re$, we observe the emergence of a single dominant frequency in the acoustic pressure signal, periodic shedding of large coherent structures, and sudden and intense heat release patterns occurring periodically [25]. Since the subsystems are co-evolving, the fluctuations in each subsystem determine as well as necessarily reflect the evolution of interactions between all subsystems. Thus, it is essential to study the role and characteristics of fluctuations during dynamical transitions.
Turbulent thermo-acoustic fluid systems are known to exhibit a transition in the acoustic field from low-amplitude aperiodic fluctuations (which is high-dimensional chaos [26, 27]) as shown in Fig. 1(d), to high-amplitude periodic oscillations (order) as shown in Fig. 1(f) with an increase in the Reynolds number ($Re$) of the inlet flow [23, 13, 28, 29, 30]. Such a transition occurs via the route of intermittency [31], which is characterized by bursts of high-amplitude periodic oscillations amidst epochs of low-amplitude chaotic oscillations in the acoustic field (Fig. 1(e)). The ordered dynamics observed at high value of $Re$ is referred as thermo-acoustic instability, characterized by large amplitude and a limit cycle attractor [23, 13].
@@
Here, we use complex networks to encode the fluctuations in the temporal dynamics of ($p^{\prime}$). A network is derived from time series of acoustic pressure fluctuations ($p^{\prime}$) obtained from experiments in a turbulent combustor, where amplitude bins are nodes, and the weight of a link connecting two nodes defines the probability of transition between the respective amplitude bins [42, 43, 44]. We refer to the resulting network as amplitude transition network. We construct these networks at distinct timescales and examine the network properties. Using network entropy, we discover that the uncertainty and the number of amplitude transitions first increase and then decrease during the transition from chaos to order in the system, delineating a non-monotonic emergence of ordered dynamics from disorder. We also characterize the amplitude scales in the fluctuations during the emergence of order using distance-based centrality measures.
The structure of the article is as follows: The details of the experimental setup and data acquisition are described in Sec. II.1. The network construction and the choice of parameters involved are explained in Sec. II.2. The measures derived from the network and their interpretations are described in Sec. III. Our conclusions are given in Sec. IV. We have also confirmed that the experimental data presented in our study show that the transition to periodic dynamics indeed occurs via intermittency, as detailed in Appendix. A.
@@ II.1 Experimental setup
We conducted experiments in a turbulent combustor with two different types of flame-holding devices. A schematic of the turbulent combustor used in this study is shown in Fig. 1(a). The turbulent combustor comprises a plenum chamber, a burner, and a combustion chamber with an extension duct. Another chamber with a relatively larger cross-sectional area, known as the decoupler, is connected to one end of the combustion duct. The decoupler reduces the loss in acoustic energy through acoustic radiation. A central shaft of diameter 16 mm through the burner supports the flame-stabilizing circular disc object, i.e., a bluff-body (Fig. 1(b)) or a vane swirler (Fig. 1(c)).
@@
Compressed air enters the plenum chamber of the combustor before passing through the burner, where the mixing of the air and gaseous fuel (liquefied petroleum gas (LPG) – 60% butane and 40% propane) takes place. The central shaft delivers fuel into the combustion chamber via four radial injection holes with diameters of 1.7 mm each. The dynamics of the unsteady acoustic pressure oscillations in the combustor is investigated as the system parameter, i.e., the Reynolds number $Re$ of the incoming flow is varied. The acoustic pressure fluctuations in the combustion chamber are acquired using a piezoelectric pressure transducer (PCB103B02, with an uncertainty of $\pm$ 0.15 Pa). The transducer is mounted at a distance of 25 mm downstream from the dump plane and records the pressure signal for 3 s at a sampling rate of 10 kHz.
@@
The $Re$ near the position of the burner is calculated as $Re=4\dot{m}D_{1}/\pi\mu D_{0}^{2}$, where $\dot{m}$ is the sum of air flow rate and fuel flow rate ($\dot{m}=\dot{m_{a}}+\dot{m_{b}}$; in g/s), $D_{0}$ is the hydraulic diameter of the burner (in mm), $D_{1}$ is the diameter of the circular bluff-body (in mm; $D_{1}=D_{0}$ for experiments with vane swirler) and $\mu$ is the dynamic viscosity of the fuel-air mixture at the experimental condition. We use the expression for dynamic viscosity $\mu$ as stated by Wilke [45] to compute the $Re$ of a binary gas mixture.
The fuel-air mixture is injected at turbulent conditions and undergoes mixing in recirculation zones created by a flame stabilizing mechanism. Here, we perform two sets of experiments using two types of flame stabilizers, namely, a bluff body and a vane swirler.
@@
The bluff body located at the end of the shaft forces the flow to turn near its shoulder, creating a large-scale recirculation zone downstream of the bluff body. On the other hand, a swirler is a non-rotating device with curved vanes (blade-like surfaces) placed radially around a central shaft, forcing the flow to turn along its vanes, creating a large-scale recirculation zone downstream of the swirler. These recirculation zones comprise hot radicals, and the large-scale recirculation vortices enhance the fuel-air mixing, which subsequently undergoes combustion. The heat released due to the combustion of the fuel-air mixture adds energy to the fluctuations in the acoustic field, leading to growth in their amplitude. Perturbations in the acoustic field influence the flame surface oscillations, leading to fluctuations in the heat release rate. Further, the acoustic and hydrodynamic fluctuations together influence the mixing and convection of the fuel-air mixture via coherent structures in the flow, which subsequently determine the dynamics of the turbulent thermo-fluid system [22, 23].
@@
For the experiments on the turbulent combustor with a bluff-body flame stabilizer, the fuel flow rate is maintained at a constant value at 1.76 g/s (grams per second), and the air flow rate is varied from 15.3 -- 29.5 g/s quasi-statically, in steps of 0.80 g/s. Hence, the corresponding $Re$ ranges from $(3.865\pm 0.031)\times 10^{4}$ to $(6.759\pm 0.054)\times 10^{4}$. Similarly, for the experiments on the turbulent combustor with swirl stabilized flame, we set a constant fuel flow rate of 0.76 g/s, and the air flow rate is varied from 6.73 -- 11.2 g/s quasi-statically, in steps of 0.20 g/s. Hence, the corresponding Reynolds number ranges from $(1.441\pm 0.011)\times 10^{4}$ to $(2.220\pm 0.018)\times 10^{4}$.
II.2 Construction of amplitude transition networks
We use the stochastic mapping network construction technique proposed by Shirazi et al. [42], which explicitly encodes the temporal sequence of the considered variable. For a discrete time series $x(t)$ = $\{x_{1},x_{2},\ldots,x_{n}\}$, we begin the network construction by grouping these data points into horizontal amplitude-bins, i.e., discrete ‘states.’
@@ II.2 Construction of amplitude transition networks
The number of bins required and the respective width of each bin are chosen such that the number of data points in each bin is approximately equal. The number of bins ($N$) needed for this formulation is given by $N\approx 2n^{2/5}$ [48].
@@ II.3 Measures from amplitude transition network
We construct amplitude transition networks using the normalized time series of acoustic pressure fluctuations ($p^{\prime}$). The normalization is done with the maximum value of the acoustic pressure during the state of thermo-acoustic instability, specifically the maximum value of $p^{\prime}$ obtained at the control parameter ($Re=(6.759\pm 0.054)\times 10^{4}$ for experiments in bluff-body combustor and $Re=(2.220\pm 0.018)\times 10^{4}$ for experiments in swirl combustor). As described in Sec. II.2, a weighted and directed network is constructed for each dynamical state at different time scales $\tau$. The topology of the constructed network encodes the temporal order of fluctuations in the dynamics. Periodic oscillations would result in a regular network topology, where connections exist in orderly fashion between specific amplitude bins. On the other hand, chaotic fluctuations would be encoded as connections between nodes corresponding to non-adjacent amplitude bins and corresponding to different amplitude scales, and the network structure becomes more disorderly.
@@
In amplitude transition network, at a specific timescale, the number of possible transitions from one amplitude scale to another is infinitely many. Yet, only a selected multitude of these transitions are realized during a certain dynamical state, reflecting the ‘complexity’ of the fluctuation dynamics at the selected timescale. Complexity here simply refers to the number of possible realizations or transitions a variable can exhibit at a fixed timescale [54, 55]. For example, a periodic signal has the least complexity as only certain amplitude transitions are realized in an orderly manner in the system. Whereas, during chaos, many amplitude transitions are realized, and the occurrence of any one of these transitions is highly uncertain. Hence, we emphasize that entropy is a measure of complexity only at the timescale at which the amplitude transitions are encoded in the network [54]. Also, Shannon entropy is equivalent to the Clausius entropy, which characterizes the macroscopic thermodynamic property of a system [56]. Thus, $\mu_{S}$ of an amplitude transition network encodes the uncertainty in temporal fluctuations of a dynamical system and also quantifies the thermodynamic disorder of the system.
@@
Furthermore, we explore the diversity in the magnitude of jumps between different amplitude scales associated with fluctuations. To do so, we use path-based centrality measures derived from amplitude transition networks. A path from one node to another in the network denotes the jump from one amplitude bin to another in the time series. Thus, we compute distance-based metrics, namely the characteristic path length (CPL) and mean betweenness centrality ($\overline{C}_{BC}$).
@@
The CPL of a network is the mean number of steps needed to reach one node from another in the network [57, 58]. For amplitude transition networks, CPL encodes the magnitude of jump in amplitude scale due to fluctuations between different amplitude bins (nodes). For example, in a periodic time series, any amplitude bin (node) is predominantly connected to specific other amplitude nodes in an orderly manner. The differences in amplitude scales are small and of similar magnitude when the transition occurs from any one bin to its neighboring bin, and this is reflected as a high value of CPL. On the other hand, a small value of CPL indicates that fluctuations occur between nodes of very different amplitude scales, such as during chaotic dynamics.
@@
The betweenness centrality of a node quantifies the extent to which a node lies in the path between other nodes of the network [59, 58]. For an amplitude transition network, the mean betweenness centrality $\overline{C}_{BC}$ quantifies the order in which transition occurs between different amplitude scales. If there are direct transitions between bins of diverse amplitude scales, such as during chaotic fluctuations, then the value of $\overline{C}_{BC}$ will be low. However, for a periodic time series, the connections are mostly between specific amplitude nodes, and any node lies ‘in between’ the path of many other node pairs with different amplitude scales, resulting in a high value for $\overline{C}_{BC}$.
@@ III Results
Figure 3 shows the variation in $\mu_{S}$ of networks derived at different transition time scales $\tau$ (in $10^{-4}$ s; as the sampling frequency of the time series is 10 kHz) from the acoustic pressure signal as the control parameter ($Re$) is varied quasi-statically in experiments. We find that the average network entropy $\mu_{S}$ varies non-monotonically with $Re$ during the transition from chaos to order in both bluff-body (Fig. 3(a)) and swirl stabilized combustors (Fig. 3(b)) in a similar manner for several transition time-scales $\tau$. We define the critical Reynolds number $Re_{c}$ as the control parameter $Re$ for which the average network entropy $\mu_{S}$ obtained from the time series of $p^{\prime}$ attains a maximum value. For the bluff-body stabilized combustor, the $Re_{c}$ is $(4.966\pm 0.040)\times 10^{4}$, whereas for the swirl combustor, it is $(1.724\pm 0.014)\times 10^{4}$. In Appendix. B, we show that the trend in network entropy is largely unaffected with variation in the other network construction parameter, i.e., the number of amplitude bins ($N$).
Figure 3: The average Shannon entropy derived from amplitude transition networks, which quantifies the amount of disorder, first increases and then decreases during the transition from chaos to order for different transition time-scales $\tau$ (in $10^{-4}$ s). (a) Variation of average network entropy ($\mu_{S}$) and root-mean-square of the non-dimensional pressure fluctuations $p^{\prime}_{rms}$ with the flow control parameter $Re$ for a bluff-body stabilized turbulent combustor. The shaded region corresponds to the region where $Re<(Re_{c})_{\text{bluff}}$ where $(Re_{c})_{\text{bluff}}=4.966\times 10^{4}$. (b) Variation of average network entropy ($\mu_{S}$) and root-mean-square of the pressure fluctuations ($p^{\prime}_{rms}$) with the flow control parameter $Re$ for a swirl stabilized turbulent combustor. The shaded region corresponds to the region where $Re<(Re_{c})_{\text{swirl}}$ where $(Re_{c})_{\text{swirl}}=1.724\times 10^{4}$.
Figure 4: (a) Variation of average Shannon entropy ($\mu_{S}$) of the network constructed from the time series corresponding to the dominant acoustic mode at different transition time scales $\tau$ (in $10^{-4}$ s). We chose the optimal time-scale to be one that maximizes the information obtained from the network, i.e., $\tau_{opt}=14\times 10^{-4}$ s for both bluff-body and swirl stabilized turbulent combustors. (b) $\mu_{S}$ from both bluff-body and swirl combustor corresponding to $\tau_{opt}$ exhibits a power law scaling $(\mu_{S})_{\text{bluff}}\propto(p^{\prime}_{rms})^{\alpha_{i}}$, and $(\mu_{S})_{\text{swirl}}\propto(p^{\prime}_{rms})^{\beta_{j}}$, respectively with similar exponents ($i$ and $j$in {1,2}). The power law exponents for a bluff-body combustor when $Re<(Re_{c})_{\text{bluff}}$ is $\alpha_{1}=0.09\pm 0.025$, and when $Re>(Re_{c})_{\text{bluff}}$ is $\alpha_{2}=-0.20\pm 0.009$. Similarly, the power law exponents for a swirl combustor when $Re<(Re_{c})_{\text{swirl}}$ is $\beta_{1}=0.15\pm 0.049$, and when $Re>(Re_{c})_{\text{swirl}}$ is $\beta_{2}=-0.21\pm 0.020$.
Figure 5: Quantifying the diversity in amplitude scale of fluctuations using distance-based network topological metrics. (a) Variation of characteristic path length (CPL) of the network constructed from time series of non-dimensional $p^{\prime}$ of bluff-body and swirl stabilized turbulent combustor with $p^{\prime}_{rms}$ (i.e., for the entire range of control parameter $Re$) at $\tau_{opt}$. The measure from both combustors exhibits the same power law scaling $\text{CPL}\propto(p^{\prime}_{rms})^{\gamma}$, with exponent $\gamma=0.28\pm 0.016$. (b) Variation of average betweenness centrality ($\overline{C}_{BC}$) of network constructed from time series of non-dimensional $p^{\prime}$ of bluff-body and swirl stabilized turbulent combustor with $p^{\prime}_{rms}$ at $\tau_{opt}$. $\overline{C}_{BC}$ from the bluff-body and swirl combustor exhibits a similar power law scaling $\overline{C}_{BC}\propto(p^{\prime}_{rms})^{\kappa}$ with exponents $\kappa=0.48\pm 0.033$ and $\kappa=0.47\pm 0.036$, respectively.
@@
Next, we chose an optimal transition timescale ($\tau_{opt}$) as the value of $\tau$ for which $\mu_{S}$ of the amplitude transition network of $p^{\prime}$ during the state of limit cycle oscillations is maximum and/or saturates. We obtain $\tau_{opt}=14\times 10^{-4}$ s for both the bluff-body and the swirl combustor (refer Fig. 4(a)), which is approximately $T/4$ s and $T/6$ s, respectively (where $T$ is the time period of $p^{\prime}$ during thermo-acoustic instability in each combustor configuration). The variation of $\mu_{S}$ with $p^{\prime}_{rms}$ at $\tau_{opt}$ is shown in Fig. 4(b). We find that $\mu_{S}$ exhibits a power law scaling with $p^{\prime}_{rms}$ and the power law exponents are similar for the results obtained from both set of experiments (bluff-body and swirler configuration). Such scaling relations are of particular interest as it helps relate the uncertainty in amplitude transitions to the amplitude itself. Evidently, this relation is similar for different combustors and indicates universality in the features of the dynamical transition rather than dependence on the specific configuration of the system.
@@
The non-monotonicity of entropy during the emergence of order reveals that during the transition from chaotic to periodic dynamics, there is an interplay between the mechanisms that promote order and disorder. A thermo-acoustic system is a complex system where the acoustic, combustion, and hydrodynamic subsystems are nonlinearly coupled. Positive feedback between these subsystems and spatially organized interactions can lead to synchronized co-evolution of variables in these subsystems, promoting pattern formation and ordered dynamics [24, 61]. On the other hand, as we increase the Reynolds number of the inlet flow, the turbulent intensity in the confined duct increases, promoting disorder and aperiodic fluctuations in the acoustic pressure. Note that both the mechanisms promoting aperiodic and periodic fluctuations in $p^{\prime}$ occur throughout all dynamical states during the transition from chaos to order. Our findings imply that disorder created by increasing turbulence intensity is prominent until a critical control parameter $Re_{c}$. Beyond $Re_{c}$, the nature of interactions between subsystems changes, perhaps becoming more spatially organized [24], and hence the periodicity is sustained for longer epochs. As a result, disorder decays monotonically despite the increase in $Re$ beyond $Re_{c}$. Moreover, the decrease in $\mu_{S}$ for $Re>Re_{c}$ is a precursor for an impending thermo-acoustic instability.